OpenStudy (anonymous):

What is the exact value of cos(2 0/) if tan 5/12 and 0/ is in quadrant 3 trigonometry 0/ is the degree?

OpenStudy (anonymous):

@iambatman

Nnesha (nnesha):

\[\cos(2^0)\] ??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

2^0 is 1

OpenStudy (anonymous):

It is the 0 with the slash in the middle of it I think its the same

OpenStudy (anonymous):

Oh, two theta

Nnesha (nnesha):

hm then its suppose to be 1\[x^0 =1 ~~ 3^0 = 1\]

Nnesha (nnesha):

lolol theta lol

OpenStudy (anonymous):

lol

Nnesha (nnesha):

rofl lol :P okay so \[\huge\rm cos(2 \theta)\] likethis lol

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

In quad 3, tan and cotan are positive everything else is negative

OpenStudy (anonymous):

The question is multiple choice too

OpenStudy (anonymous):

25/169 -287/169 -119/169 119/169

Nnesha (nnesha):

 x and y both are negative